# Strategic Pricing and Ranking in Recommendation Systems with Seller Competition

### v1

@@ Abstract
The core contribution of this paper lies in characterizing the equilibrium structure of the limit game. We show that when sellers are of different strengths, the standard Nash equilibrium does not exist due to discontinuities in utilities. We instead establish the existence of a novel equilibrium solution, namely $\mu$-connected equilibrium cycle’ ($\mu$-EC), which captures oscillatory strategic responses at the equilibrium. Unlike the (pure) Nash equilibrium, which defines a fixed point of mutual best responses, this is a set-valued solution concept of connected components. This novel equilibrium concept identifies a Cartesian product set of connected action profiles in the continuous action space that satisfies four important properties: stability against external deviations, no external chains, instability against internal deviations, and minimality. We extend a recently introduced solution concept equilibrium cycle to include stability against measure-zero violations and, by avoiding topological difficulties to propose $\mu$-EC.
@@ 1 Introduction
While classical models focus on optimizing rankings or prices in isolation (e.g., [1, 2]), modern digital marketplaces face a more intricate challenge: how to jointly align pricing, ranking, seller incentives, and customer behaviour, especially in settings where the sellers actively respond to and attempt to influence the platform policies. This interaction leads to a two-sided strategic tension—between platforms optimizing revenue and sellers seeking favourable positioning, which remains under-explored in the recommendation systems literature.
The paper investigates this tension by modelling the interaction between sellers and a revenue-maximizing platform. We focus on a commission-based design, where sellers offer monetary incentives (commissions) to the platform in the hope of securing higher visibility. The platform, in response, constructs a personalized ranked menu and depending on the setting, may also determine the item prices. It optimizes its revenue by accounting for both the offered commissions and customer click-through behaviour.
This interplay is modelled using a Stackelberg game framework with sellers as leaders competing for customer attention by strategically choosing commissions of the platform. We consider two distinct scenarios:
@@
These models capture a wide spectrum of real-world platforms—from those retaining full control over prices to marketplaces where prices are fixed externally (e.g., MSRP or multi-platform listings). Unlike existing literature that often decouples pricing and ranking decisions, our model integrates these choices. We uncover a rich feedback loop: seller commission strategies influence menu placement, which in turn affects the customer choices and the platform revenue, thereby feeding back into the seller’s revenue and hence the strategic considerations. This strategic interdependence gives rise to novel equilibrium behaviour not captured by classical solution concepts. Our work makes following key contributions:
@@
    Computational framework for platform optimization: We pose the platform’s response as an MDP (Markov Decision Process) and solve it via backward induction. When the customer drop-off rate is high, we derive the closed-form solutions for platform optimization problem, reducing the complexity from exponential to linear in menu size—thus making our model scalable.
@@
    Equilibrium characterization via $\mu$-connected equilibrium cycle: Our central theoretical innovation is the introduction of a novel set-valued solution concept: the $\mu$-connected Equilibrium Cycle ($\mu$-EC). In games with discontinuities, caused here by rank jumps in menus, the classical pure Nash equilibria may fail to exist. The $\mu$-EC captures stable yet cyclic equilibrium: sellers perpetually adjust their commissions in reaction to competitors, but the strategy profiles remain confined within a minimal, recurrent set. We establish the existence of $\mu$-ECs as a solution for the scenarios with asymmetric sellers. Interestingly, the weaker sellers act as a threat that constrains the top-ranked sellers to offer commissions higher than the weak sellers’ potential.
This cyclical strategic behaviour mirrors phenomena observed empirically in auction and advertising systems (e.g., [3]), where sellers repeatedly adjust bids without convergence. Our $\mu$-EC formalism captures and generalizes this behaviour in the context of recommendation systems with strategic sellers.
This paper is structured as follows: after a brief literature review, Section 2 presents the system model. Section 3 discusses the platform’s optimization in the OP game via an MDP framework. Section 4 provides a small-$\gamma$ approximation for the OP game, while Section 5 does so for the OS game. Section 7 defines the solution concept $\mu$-EC.The paper concludes in Section 8.
@@
We adopt the cascade click model introduced by [1] and extended in [2] to capture sequential customer interactions with ranked menus. These models, while foundational for customer behaviour modelling, do not incorporate platform-driven pricing or strategic seller behaviour. In contrast, our work embeds these behavioural dynamics within a game-theoretic framework involving commission competition and platform optimization. Broadly, recommendation systems have been studied through algorithmic, behavioral, and learning-theoretic lenses, often focusing on optimizing rankings, prices, or assortments based on buyer preferences. However, such models typically treat sellers as passive entities and overlook their potential to strategically influence platform decisions. Our work departs from this by treating sellers as active agents who compete via commissions, directly affecting both ranking and pricing outcomes.
Several prior works consider platform-seller interaction using Stackelberg game formulations. Notably, [4] model a setting where sellers lead and the platform follows, but their formulation considers only pricing. We generalize this by introducing two coupled games: the OP game, where the platform jointly decides rankings and prices, and the OS game, where sellers pre-fix prices and compete over rankings. This dual formulation captures a richer set of real-world scenarios. In a different direction, [5] considers a setting where sellers are partitioned into relevance-based classes, and visibility is determined accordingly. Sellers in their model compete via pricing, but inter-class competition is restricted. Our model moves beyond such segmentation and allows for full-scale cross-seller competition, enabling all sellers to influence the outcomes through aggressive commission strategies. This interplay becomes central in our equilibrium analysis, where weaker sellers can act as credible threats to stronger ones.
Recent studies, such as [6, 7] explore strategic platform design in environments where sellers are algorithmically sophisticated and capable of deploying pricing strategies through automated tools. This enhances the significance of our analysis of strategic sellers. Additionally, there is growing attention on the dual role of platforms as intermediaries and market participants, raising concerns about fairness and market design. For example, [8] and [9] examine how platforms entering retail markets can bias outcomes through preferential control over product visibility. While the platform in our model does not directly compete with sellers, it still has a significant control through pricing and ranking decisions. It also chooses the sellers fairly among the sellers with optimal utility.
Finally, the role of commissions has also been examined in coordinated settings. Authors in [10] study a system with two sellers and a platform, showing that aligning commissions with service levels can improve collective efficiency. Unlike their cooperative framework, our model is fully non-cooperative, with sellers strategically competing over visibility through commissions. A central challenge we address is the lack of classical NE due to discontinuities in seller utilities. To resolve this, we introduce the notion of a $\mu$-EC, which captures persistent, non-convergent but structurally stable oscillations in seller strategies; an empirically observed phenomenon in strategic bidding environments such as search advertising ([3]).
In summary, our work builds on and extends existing literature by modeling strategic sellers, platform-side pricing and ranking, and a non-classical equilibrium structure that reflects the complex, dynamic nature of real-world platform markets.
@@ 2.4 Utilities of different agents
while that of the sellers are $(1-\delta_{a})p_{a}\beta_{a}(p_{a})$ and $(1-\delta_{b})p_{b}(1-\beta_{a}(p_{a}))\gamma\beta_{b}(p_{b})$ respectively. These seller revenues do not include the respective production costs. We now describe the utility functions of the sellers, including the costs.
The profit of seller $a$ upon selling its product is given by $(1-\delta_{a})p_{a}-c_{m_{a}}$, where $c_{m_{a}}$ is the production cost. However, the opportunity to sell its product depends on $(\pi,{\bf p})$, and the combination of these two factors defines its overall utility. For given $({\bm{\delta}},\pi,{\bf p})$, the utility of seller $a$ is:
${U}_{a}({\bm{\delta}},\pi,{\bf p})=\sum_{k=1}^{M}\left(\mathds{1}_{\left\{a=\pi(k)\right\}}((1-\delta_{a})p^{*}_{\pi,a}-c_{m_{a}})\beta_{a}(p^{*}_{\pi,a})\prod_{j<k}\left(1-\beta_{\pi(j)}(p^{*}_{\pi,\pi(j)}\right)\gamma\right)\mbox{ for all }a\in\mathcal{I}.$ (2)
In above, when $\mathds{1}_{\left\{a=\pi(k)\right\}}$ is one, it implies that the seller is placed in position $k$ in the menu $\pi$. Next, we formally define the two games.
@@ 2.5.1 Order and price (OP) game
This policy ${\bf z}^{*}({\bm{\delta}})$ is actually derived in Section 3; for now, we assume its existence and proceed towards providing the complete description of the Stackelberg (OP) game. At the top, we have a strategic form game between $N$ sellers. Each seller chooses a commission $\delta_{a}\in[0,1]$, and the response of the follower platform, ${\bf z}^{*}({\bm{\delta}})$, influences their utilities (see (2)):
${U}_{a}({\bm{\delta}}):=\sum_{\pi}\Pi^{*}({\bm{\delta}})(\pi)\sum_{k=1}^{M}\left(\mathds{1}_{\left\{a=\pi(k)\right\}}((1-\delta_{a})p^{*}_{\pi,a}-c_{m_{a}})\beta_{a}(p^{*}_{\pi,a})\prod_{j<k}\left(1-\beta_{\pi(j)}(p^{*}_{\pi,\pi(j)}\right)\gamma\right)\forall\ a\in\mathcal{I}.$ (4)
@@ 2.5.2 OS game: order at set-prices
This price $p^{*}_{a}$ is fixed and remains unchanged during competition. The sellers then strategically set $\delta_{a}\in[0,1]$ to maximize their utility, anticipating the platform’s menu optimization.
@@
As mentioned above, one can obtain the unique ‘fair’ optimizer now in terms (only) of random menus $\Pi^{*}({\bm{\delta}})$. The utility of seller $a$ equals (see (4) and also observe that the prices are set a priori):
@@ 3.1 MDP formulation
    Finally, we have the existence of solution for (3) and the MDP for any $\bm{\delta}$}:
    $\max_{{\bar{\bf z}}=(\pi,{\bf p})}R(\bar{\bf z},\bar{\bf{x}})\stackrel{{\scriptstyle a}}{{=}}\max_{\mu,\mbox{ deterministic}}J(\mu,\bar{\bf{x}}).$
@@ 3.2 Dynamic programming equations and the optimal-fair policy
Clearly, the domain of optimization for each $(t,{\bf{x}})$ is compact and the instantaneous reward function $p\mapsto r({\bf{x}},(a,p))$ is continuous, for each $({\bf{x}},a)$. Hence, by [21, Chapter 4], there exists a solution to the DP equation (9). Consequently, there exists a solution among the deterministic policies, which serves as the second proof for the existence of a solution to (3), this time using MDP theory (recall that existence was alredy established in Lemma 1). However, more interestingly, the existence of solution for DP equation (9) also paves way for choosing an unique ‘fair’ policy among all possible optimal policy(ies) as explained in the immediate next.
@@
The above set of optimizers need not be unique for every $(t,{\bf{x}})$, however, the following ‘fair’ randomized policy is one among the optimizers, as established in [21, Chapter 4]:
@@
This policy ensures that equal preference is given to each of the ‘equal’ sellers at any $(t,{\bf{x}})$. We now formally summarize the optimality of the fair-policy below:
@@ 6.1 Accuracy of small-$\gamma$ approximation
Proof.
In view of the above result, for the exponential response function (20), it is clear that the exact optimal policy of the platform coincides with small-$\gamma$ approximate policy for all $\gamma$ for the OS game. Consequently, the utility of any seller in the OS game (7) exactly equals that in (22) and hence the approximate unified seller game coincides with the OS game for all $\gamma$ (recall in OS game, the prices are pre-set according to (5)). We now examine numerically the accuracy of the small $\gamma$ approximation for OP game.
((a))
((b))
@@
Figure 2 compares the seller utilities under the OP game and the approximate policy for varying values of $\gamma$. For $\gamma=0.2$, the two curves align almost perfectly, indicating almost exact utility matching. As $\gamma$ increases to $0.4$ and $0.7$, the curves start to diverge slightly. To better visualize these deviations, Figure 2(b) presents a zoomed-in version of the original plot shown in Figure 2(a), focusing on the regions of $\bm{\delta}$} where the differences are more noticeable: the percentage error is $1.17\%$, $2.54\%$ and $5.14\%$ respectively for $\gamma=0.2$, $\gamma=0.4$ and $\gamma=0.7$. This demonstrates that the approximate policy provides a sufficiently accurate estimates of seller utilities even for moderate values of $\gamma$. Next, we move to the analysis of the unified seller game.

### v2

@@ Abstract
The core contribution of this paper lies in characterizing the equilibrium structure of the limit game. We show that when sellers are of different strengths, the standard Nash equilibrium does not exist due to discontinuities in utilities. We instead establish the existence of a novel equilibrium solution, namely $\mu$-connected equilibrium cycle’ ($\mu$-EC), which captures oscillatory strategic responses at the equilibrium. Unlike the (pure) Nash equilibrium, which defines a fixed point of mutual best responses, this is a set-valued solution concept of connected components. This novel equilibrium concept identifies a Cartesian product set of connected action profiles in the continuous action space that satisfies four important properties: stability against external deviations, no external chains, instability against internal deviations, and minimality. We extend a recently introduced solution concept equilibrium cycle to include stability against measure-zero violations and avoid some topological difficulties to propose $\mu$-EC.
@@ 1 Introduction
While classical models focus on optimizing rankings or prices in isolation (e.g., [1, 2]), modern digital marketplaces face a more intricate challenge: how to jointly align pricing, ranking, seller incentives, and customer behaviour, especially in settings where the sellers actively respond to and attempt to influence the platform policies. This interaction leads to a two-sided strategic tension—between platforms optimizing revenue and sellers seeking favorable positioning, which remains under-explored in the recommendation systems literature.
The paper investigates this tension by modelling the interaction between sellers and a revenue-maximizing platform. We focus on a commission-based design, where sellers offer monetary incentives (commissions) to the platform in the hope of securing higher visibility. The platform, in response, constructs a personalized ranked menu and depending on the setting, may also determine the item prices. It optimizes its revenue by accounting for both the offered commissions and customer click-through behaviour (e.g., as in [1, 2]).
This interplay is modelled using a Stackelberg game framework with sellers as leaders competing for customer attention by strategically choosing commissions for the platform. We consider two distinct scenarios:
@@
These models capture a wide spectrum of real-world platforms—from those retaining full control over prices to marketplaces where prices are fixed externally (e.g., MSRP or multi-platform listings). In the current platform-aided marketplaces, the sellers set the prices, however the prices are significantly influenced by the norms and the policies of platform (e.g., see [3, 4] on how major platforms like Walmart, Flipkart influence seller prices). This motivates us to consider the above two games and analyse them individually as well as to obtain a comparative study. Further, unlike the existing literature that often decouples pricing and ranking decisions, our model integrates these choices. We thus uncover a rich feedback loop: seller commission strategies influence menu placement, which in turn affects the customer choices and the platform revenue, thereby feeding back into the seller’s revenue and hence the strategic considerations. This strategic interdependence gives rise to novel equilibrium behaviour not captured by classical solution concepts. Our work makes the following key contributions:
@@
    Computational framework for platform optimization: We pose the platform’s response as an MDP (Markov Decision Process) and solve it via backward induction. When the customer drop-off rate is high, we derive the closed-form solutions for platform optimization problem, thereby reducing the complexity from exponential to linear in menu size, and thus making our model scalable.
@@
    Equilibrium characterization via $\mu$-connected equilibrium cycle: Our central theoretical innovation is the introduction of a novel set-valued solution concept: the $\mu$-connected Equilibrium Cycle ($\mu$-EC). In games with discontinuities, caused here by rank jumps in menus, the classical pure Nash equilibria does not exist under certain scenarios. The $\mu$-EC captures stable yet cyclic equilibrium for such case-studies: sellers perpetually adjust their commissions in reaction to competitors, but the strategy profiles remain confined within a minimal, recurrent set. We establish the existence of $\mu$-ECs as a solution for the scenarios with asymmetric sellers. Interestingly, the weaker sellers act as a threat that constrains the top-ranked sellers to offer commissions higher than the weak sellers’ potential.
    This kind of cyclical strategic behaviour mirrors phenomena observed empirically in auction and advertising systems (e.g., [5]), where sellers repeatedly adjust bids without convergence. Our $\mu$-EC formalism captures and generalizes this behaviour in the context of recommendation systems with strategic sellers.
  • 6.
    The new notion $\mu$-EC is a generalization of the recently proposed solution concept, namely Equilibrium cycle ([6]): the $\mu$-EC allows measure-zero exceptions and avoids some topological difficulties. It identifies a connected set of action profiles that are (i) stable against external deviations, (ii) contain no external chains, (iii) unstable against internal deviations, and (iv) are minimal. The key refinement lies in permitting temporary deviations that leave the set but quickly return, ensuring that long-run dynamics remain confined within the $\mu$-EC.
@@
We adopt the cascade click model introduced by [1] and extended in [2, 7, 8, 9] to capture customer interactions with ranked menus. These models, while foundational for customer behaviour modelling, do not incorporate platform-driven pricing or strategic seller behaviour. In contrast, our work embeds these behavioural dynamics within a game-theoretic framework involving commission competition and platform optimization. Broadly, recommendation systems have been studied through algorithmic, behavioral, and learning-theoretic lenses, often focusing on optimizing rankings, prices, or assortments based on buyer preferences. However, such models typically treat sellers as passive entities and overlook their potential to strategically influence platform decisions. Our work departs from this by treating sellers as active agents who compete via commissions, directly affecting both ranking and pricing outcomes.
On the other hand, authors in ([10, 11, 12]) have studied platform-seller strategic interactions using Stackelberg game formulations, where sellers are the suppliers and the platform is a retailer. These strands of literature do not consider the design of menus that strategically present the products of multiple sellers to customers in an optimal order. By contrast, we consider both pricing and ranking decisions in two distinct (OS and OP) games.
Thus, we have pieces of work that focus only on ranking decisions (or optimal menu design) and other strands that consider pricing games alone. But to the best of our knowledge, we did not come across any literature that captures the framework of joint decisions, which is predominant in real-world applications. Our work, therefore, fills this gap. At the same time, it is loosely related to several other themes of work reported in the literature, a few of which we discuss below.
In the recent past, there has been a significant body of work that discusses the aggressive dynamic prices set by sellers using automated algorithms. The use of algorithms in place of human decision-making increases price fluctuations, as demonstrated in [13, 14, 15, 16]. For example, [13] states that even in 2015 about one-third of Amazon’s top-selling products were already priced using algorithms, with algorithms changing prices hundreds of times more often than human sellers. Authors in [14] show that platforms can influence these outcomes through their ranking rules and by rewarding price cuts. Overall, these studies show that platforms are not just neutral intermediaries, but they actively shape the competition through their decisions. This line of work confirms that sellers also behave strategically, providing strong motivation for our framework which explicitly incorporates strategic sellers into recommendation systems.
Another related branch of literature [17, 18] discusses platforms that directly compete with sellers on certain products, raising fairness concerns. While the platform in our model does not directly compete with sellers for market share, our framework can easily be extended to study such scenarios by setting commission for in-house products to 100 %, and allowing the remaining sellers to compete as before.
The role of commissions in platform-based markets has also been studied extensively in the literature, for example, [19, 20, 21]. Authors in [20] analyze a cooperative setting with two sellers and a platform, where pricing, service quality, and commissions are jointly determined. In their framework, the platform benefits from intense seller competition, much like in our model, while the sellers prefer cooperation to increase their profits. However, these pieces of work do not consider the influence of the ranking decisions. Further, unlike such cooperative scenarios, our framework is fully non-cooperative: sellers choose commissions purely to maximize their own revenue, and the platform optimizes its policy in response. This non-cooperative setting introduces new challenges. In particular, we encounter situations where a classic pure-strategy Nash equilibrium does not exist due to discontinuities in seller’s payoff. As already mentioned, to address this, we introduce the concept of $\mu$-EC, that can capture cyclic strategic outcomes.
This work also refines the recently introduced cyclic equilibrium outcome, namely Equilibrium cycle introduced in [6], by proposing a refined $\mu$-EC. In this regard, we now list below strategic cyclic phenomenon observed in other applications. In related empirical contexts such as airline and gasoline markets, researchers have documented recurring price wars known as Edgeworth cycles. These cycles describe a pattern where firms repeatedly undercut each other until prices fall close to marginal cost, after which prices reset to higher levels and the cycle restarts [22, 23, 24]. Similar dynamics also have been observed in online advertising, where the competing bidders adjust bids in persistent cycles without convergence [5]. The persistence of such dynamics highlights the limitations of classical equilibrium notion of Nash equilibrium in explaining real-world competitive outcomes. Our $\mu$-EC framework offers a natural theoretical explanation for these phenomena by formally capturing oscillatory yet stable strategy profiles.
In summary, our work brings together insights from cascade models of consumer behavior, pricing and ranking optimization, seller–platform competition, algorithmic pricing, and equilibrium analysis in discontinuous games.
This paper is structured as follows: Section 2 presents the system model. Section 3 discusses the platform’s optimization via an MDP framework. We discuss small-$\gamma$ approximation in Sections 4 and 5. Section 7 defines the solution concept $\mu$-EC. We conclude in Section 8.
@@ 2.4 Utilities of different agents
while that of the sellers are $(1-\delta_{a})p_{a}\beta_{a}(p_{a})$ and $(1-\delta_{b})p_{b}(1-\beta_{a}(p_{a}))\gamma\beta_{b}(p_{b})$ respectively. These seller revenues do not include the respective production costs. We now describe the utility functions of the sellers, including the costs. The profit of seller $a$ upon selling its product is given by $(1-\delta_{a})p_{a}-c_{m_{a}}$, where $c_{m_{a}}$ is the production cost. However, the opportunity to sell its product depends on $(\pi,{\bf p})$, and the combination of these two factors defines its overall utility. For given $({\bm{\delta}},\pi,{\bf p})$, the utility of seller $\color[rgb]{0,0,0}{a\in\mathcal{I}}$} equals:
${U}_{a}({\bm{\delta}},\pi,{\bf p})=((1-\delta_{a})p^{*}_{\pi,a}-c_{m_{a}})\beta_{a}(p^{*}_{\pi,a})\prod_{j<k}\left(1-\beta_{\pi(j)}(p^{*}_{\pi,\pi(j)}\right)\gamma,\quad\mbox{ if }a=\pi(k)\mbox{ for some }k\leq M,$ (2)
i.e., if placed in $k$-th position with $k\leq M$, otherwise ${U}_{a}({\bm{\delta}},\pi,{\bf p})=0$. Next, we formally define the two games.
@@ 2.5.1 Order and price (OP) game
This policy ${\bf z}^{*}({\bm{\delta}})$ is actually derived in Section 3; for now, we assume its existence and proceed towards providing the complete description of the Stackelberg (OP) game. At the top, we have a strategic form game between $N$ sellers. Each seller chooses a commission $\delta_{a}\in[0,1]$, and the response of the follower platform, ${\bf z}^{*}({\bm{\delta}})$, influences their utilities (see (2), $\mathds{1}_{\left\{a=\pi(k)\right\}}$ is indicator of the seller being placed in $k$-th position in $\pi$):
${U}_{a}({\bm{\delta}}):=\sum_{\pi}\Pi^{*}({\bm{\delta}})(\pi)\sum_{k=1}^{M}\left(\mathds{1}_{\left\{a=\pi(k)\right\}}((1-\delta_{a})p^{*}_{\pi,a}-c_{m_{a}})\beta_{a}(p^{*}_{\pi,a})\prod_{j<k}\left(1-\beta_{\pi(j)}(p^{*}_{\pi,\pi(j)}\right)\gamma\right)\forall\ {\color[rgb]{0,0,0}{a\in\mathcal{I}}}.$ (4)
@@ 2.5.2 OS game: order at set-prices
This price $p^{*}_{a}$ is fixed and remains unchanged during competition. Sellers then strategically set $\delta_{a}\in[0,1]$ to maximize their utility, anticipating the platform’s menu optimization.
@@
As mentioned above, one can obtain the unique ‘fair’ optimizer now in terms (only) of random menus $\Pi^{*}({\bm{\delta}})$. The utility of seller $\color[rgb]{0,0,0}{a\in\mathcal{I}}$} equals (see (4) and also observe that the prices are set a priori):
@@ 3.1 MDP formulation
    Finally, we have the existence of solution for (3) and the MDP for any $\bm{\delta}$}: max_¯z = (π, p) R(¯z, ¯x) =a max_μ, deterministicJ( μ, ¯x) .
@@ 3.2 Dynamic programming equations and the optimal-fair policy
Clearly, the domain of optimization for each $(t,{\bf{x}})$ is compact and the instantaneous reward function $p\mapsto r({\bf{x}},(a,p))$ is continuous, for each $({\bf{x}},a)$. Hence, by [34, Chapter 4], there exists a solution to the DP equation (9). Consequently, there exists a solution among the deterministic policies, which serves as the second proof for the existence of a solution to (3), this time using MDP theory (recall that existence was alredy established in Lemma 1). However, more interestingly, the existence of solution for DP equation (9) also paves way for choosing an unique ‘fair’ policy among all possible optimal policy(ies) as explained in the immediate next.
@@
The above set of optimizers need not be unique for every $(t,{\bf{x}})$, however, the following ‘fair’ randomized policy is one among the optimizers, as established in [34, Chapter 4]:
@@
This policy ensures that equal preference is given to each of the ‘equal’ sellers at any $(t,{\bf{x}})$, which takes care of fairness. The problem of fair allocations dates back to several decades ago: for example, in the context of wireless allocations, authors in [35] have discussed about proportional fair allocation of channels to various competing users. There were several other papers, hence after, that discuss fair allocations. Importance is given to such allocations, as these directly reflect the quality of service (QoS), which in turn reflects the customer satisfaction. In our context, fair-allocations imply ordering the ‘equal’ sellers in an ‘equal’ manner. When the platform implements (11), on average the visibility of all the ‘equal’ sellers would be the same. Once again such allocations reflect better QoS to sellers, which in turn can improve the credibility of the platform (thereby improving the interest across sellers).
We now formally summarize the optimality of the fair-policy (11) below:
@@ 6.1 Accuracy of small-$\gamma$ approximation
In view of the above result (proof in C, for the response function (20), it is clear that the exact optimal policy of the platform coincides with small-$\gamma$ approximate policy for all $\gamma$ for the OS game. Consequently, the utility of any seller in the OS game (7) exactly equals that in (22) and hence the approximate unified seller game coincides with the OS game for all $\gamma$ (recall in OS game, the prices are pre-set according to (5)). We now examine numerically the accuracy of the small $\gamma$ approximation for OP game.
(a) Seller utility under OP and approximate policy for different values of $\gamma$
(b) Zoomed-in view of Figure 2a highlighting regions with visible deviations between the curves
@@
Figure 2 compares the seller utilities under the OP game and the approximate policy for varying values of $\gamma$. For $\gamma=0.2$, the two curves align almost perfectly, indicating almost exact utility matching. As $\gamma$ increases to $0.4$ and $0.7$, the curves start to diverge slightly. To better visualize these deviations, Figure 2b presents a zoomed-in version of the original plot shown in Figure 2a, focusing on the regions of $\bm{\delta}$} where the differences are more noticeable: the percentage error is $1.17\%$, $2.54\%$ and $5.14\%$ respectively for $\gamma=0.2$, $\gamma=0.4$ and $\gamma=0.7$. This demonstrates that the approximate policy provides a sufficiently accurate estimates of seller utilities even for moderate values of $\gamma$. Next, we move to the analysis of the unified seller game.